is the points A(2,1),B(-4,1),C(-1,-3) Collinear or noncollinear
Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(75887)

If the slopes from point to point are all the same the points are collinear.
-----------------
is the points A(2,1),B(-4,1),C(-1,-3) Collinear or noncollinear
---------------
from A to B: slope = 0
from B to C; slope = (-3-1)/(-1--4) = -4/3
----------------
slopes are not the same so the points are not collinear
===============
Cheers,
Stan H.

Answer by Fombitz(32388)

For points to be collinear, the change in y divided by the change in x is constant and equal to the slope of the line connecting the two points.
The slope between A and B is zero, since the change in y is zero (1-1=0).
The slope between A and C is not zero, since the change in y is not zero (-3-1=-4).
Therefore the points are not collinear.